40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3713     Entry time: Thu Oct 14 01:09:17 2010
Author: yuta 
Type: Update 
Category: IOO 
Subject: tried to lock MC but failed 

(Rana, Koji, Kiwamu, Suresh, Yuta)

 We attempted to lock the MC and finally got flashes of the MC, but no luck. 

Tomorrow we are going to check every components one by one to make sure if everything is okay or not.


Background:
 MC suspensions are well damped now.
 We need MC locked for the alignment of the in-vac optics.

Issues:

  These are the issues which we are going to fix.
1. DC alignment of the MC2 suspension doesn't seem to be working correctly. (see here)

  We should check the satellite box and the cable connection.

  The coils look like woking fine because we can kick MC2 by using each of the coils.


2. Incomplete modematching.

The spot size of the reflected light from MC1 looked like bigger. 

3. beam axis of the injection light to MC1

ELOG V3.1.3-